OpenStudy (anonymous):

ratio 1:40 as a percent???

OpenStudy (anonymous):

1:40=1/40 percentage you need to multiply the ratio by 100. 1/40*100=100/40=2.5%

OpenStudy (anonymous):

thank you!!!!!!